OpenStudy (anonymous):

Discuss possible conditions on a and b that would make this a valid equation: (a-b)^2 does not equal a^2-b^2.

OpenStudy (anonymous):

First of all, evaluate (a-b)^2.

OpenStudy (anonymous):

a^2-2ab+b^2

OpenStudy (anonymous):

Okay now if we say: a^2-b^2 = a^2-2ab+b^2 Subtract a^2 and get: -b^2 = -2ab+b^2 Subtract b^2 and get: -2b^2 = -2ab Divide by -2 b^2 = ab Divide by b: a=b

OpenStudy (anonymous):

So we have two cases, a=b and b=0, where they are equal. If this isn't true, then (a-b)^2 does not equal a^2-b^2.